40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5473     Entry time: Tue Sep 20 02:21:10 2011
Author: Koji 
Type: Update 
Category: LSC 
Subject: LSC MEDM screen cleaning up 

I have made some cleaning up of the LSC-related MEDM screens.

- LSC overview screen: ADC OVFL and WFAA indicators are now correctly matched to it associated PD signals.

- Whitening screens now have the correct indication of the associated PD signals.

- LSC Ctrl screen, which is invoked from the overview screen by clicking the servo filters, now has the switches of the servo filters.

- LSC tab of the sitemap was cleaned up by removing the broken links.

ELOG V3.1.3-